# Full spectrum of Love numbers of Reissner-Nordström black hole in $D$-dimensions

### v1

@@ 1 Introduction
Tidal Love numbers describe the deformability of a black hole under the influence of external force. It has been demonstrated that conservative linear tidal Love numbers vanish for several families of asymptotically flat black holes in four dimensions, including Schwarzschild [1], Kerr [2, 3, 4] and Reissner-Nordström (RN) [5, 6, 7] black holes. It was also found that for Kerr black holes, the conservative tidal Love numbers remain vanishing at nonlinear level [8, 9, 10, 11]. These results provide strong support to the no-hair theorem, showing that black holes in four-dimensional general relativity are remarkably rigid objects. On the other hand, Love numbers are generically nonvanishing for charged black holes coupled to charged scalar or fermion [12, 13, 14, 15] and black holes in higher dimensions [16, 17, 18, 19], in modified gravitational theories [20, 21, 22, 23, 24, 25], in the presence of a cosmological constant [26, 27], or in astrophysical environments [28, 29, 30, 31, 32]. Its potential observability by future gravitational wave experiments makes Love numbers useful theoretical diagnostics.
@@
When restricted to $D=4$, our master equations match with results in [42, 43, 44]. In $D>4$, the metric perturbations contain also an intrinsic tensor mode that propagates on its own without mixing with others. This constitutes the simplest perturbation channel. The other components of the metric perturbations inevitably mix with the electromagnetic counterparts. Due to the different methods used in diagonalizing the perturbations, our master equations do not look the same as those in [6]. However, the Love numbers in the tensor and vector sectors do agree with those obtained in [6]. In particular, we confirm that all the Love number vanishes for $4D$ RN black holes [6, 7], a feature similar to the Schwarzschild black hole [1, 2, 39]. In $D>4$, we also recover earlier results for the tensor and vector Love numbers [6]. So far, the scalar type Love numbers of RN black hole have not been studied. We recall that in the case of Schwarzschild black hole, when $\tilde{\ell}$ is integer valued, all scalar type Love numbers vanish. Hence, it should be interesting to see if similar features appear also in RN black hole. Our results show that this is indeed the case, which may imply the existence of certain accidental symmetry, as was found for other cases, including $4D$ Schwarzschild black hole [45, 46, 47] and RN black hole [7] and certain higher dimensional black branes [48] or black holes [49].
@@ 3.1 Tensor Love numbers in diverse dimensions
    In this way of parameterization, $\Phi_{\rm resp}$ and $\Phi_{\rm tidal}$ are analytic functions in a neighborhood of $z=0$. Demanding regularity at the horizon, the ratio between $A$ and $B$ can be fixed and the solution takes the form
@@
    The numerical value of $k_{T}$ is determined by matching the solutions near the horizon $(z=1)$ and at infinity $(z=0)$. In Fig.1, we present both the analytical (solid line) and numerical (points) results for the tensor Love numbers at $\tilde{\ell}=1/3$. The excellent agreement between the numerical and theoretical values demonstrates the validity of our method.
@@ 3.2 Effective action of vector type perturbations
Unlike the tensor sector (30), the Regge-Wheeler mode in the gravitational perturbation and the $s=1$ mode in the Maxwell field are both vector-type perturbations, which couple with each other. As a result, the action for vector sector is more complex than that for tensor sector (30). In this section, we first introduce an auxiliary field to eliminate the non-physical degrees of freedom, thereby obtaining the two-dimensional gravitational-Maxwell coupling action. Then, we decouple the action by a specific linear combination.
Since both the gravitational and Maxwell field perturbations contain vector modes, they are coupled to each other. As a result, the effective Lagrangian of this sector is of the form
@@
In the $s=1$ sector, the gauge-invariant Love numbers are encoded in the magnetic components of the Maxwell field strength and linearized Weyl tensor,
@@ 3.3 Vector-type Love numbers in diverse dimensions
Similarly, through a coordinate transformation $x\rightarrow z=\left(\frac{r_{+}}{r}\right)^{D-3}=1/((1-\sigma)x+1)$, we can convert (65) into the standard Fuchsian form. For equations of this type, the indices near the singular point $z=0$ are $\tilde{\ell}$ and $\tilde{\ell}+1$ and the analysis can be divided into the two cases:
@@
    Near spatial infinity ($z\rightarrow 0$) , the asymptotic behaviors of $\Psi_{RW}$ and $\Psi_{V}$ is then
    $r^{\frac{2-D}{2}}\Psi_{RW}|_{z\rightarrow 0}$ $=z^{\tilde{\ell}+1}\left(\alpha_{1}B_{+}R_{+}+\alpha_{2}B_{-}R_{-}\right)\ln z+z^{-\tilde{\ell}}(\alpha_{1}B_{+}+\alpha_{2}B_{-})\ ,$ (73)
@@
    $k_{RW}=\frac{\alpha_{1}B_{+}R_{+}+\alpha_{2}B_{-}R_{-}}{\alpha_{1}B_{+}+\alpha_{2}B_{-}}\ln{z}\ ,\quad k_{V}=\frac{\alpha_{3}B_{+}R_{+}+\alpha_{4}B_{-}R_{-}}{\alpha_{3}B_{+}+\alpha_{4}B_{-}}\ln{z}\ ,$ (75)
    which clearly exhibits gravito-photon mixing. Thus in general, the Love number depends on the ratio of the amplitudes on the horizon. Following [6], when computing the Love number for a specific mode, we set the tidal field of the other mode to zero, thereby fixing the relative amplitude between the two modes. Thus, when calculating $k_{RW}$ ($k_{V}$), we set the coefficient of the $z^{-\tilde{\ell}}$ term in $\Psi_{V}|_{z\rightarrow 0}$ ($\Psi_{RW}|_{z\rightarrow 0}$) to zero, fixing the ratio of the amplitudes to be $B_{+}/B_{-}=-\alpha_{4}/\alpha_{3}$ ($B_{+}/B_{-}=-\alpha_{2}/\alpha_{1}$). This allows us to derive $k_{RW}$ and $k_{V}$ analytically. It can be verified that in $D=4$, $R_{\pm}$ vanishes, resulting in a tidal field $\Phi_{\rm tidal\pm}$ given by polynomials of degree less than $2\tilde{\ell}+1$ and thus $k_{RW}=k_{V}=0$. For $D=5$, we list values of the Love numbers (logarithm omitted) for various $\tilde{\ell}$ in Fig.2.
    Figure 2: These plots show how vector type Love numbers in $D=5$ behave as we change $q/\mu$ and $L$. $\ln z$ is omitted in the plots.
@@
    $k_{RW}=\frac{\alpha_{2}\alpha_{3}k_{-}-\alpha_{1}\alpha_{4}k_{+}}{\alpha_{2}\alpha_{3}-\alpha_{1}\alpha_{4}}\ ,\quad k_{V}=\frac{\alpha_{1}\alpha_{4}k_{-}-\alpha_{2}\alpha_{3}k_{+}}{\alpha_{1}\alpha_{4}-\alpha_{2}\alpha_{3}}\ .$ (77)
    For given $D$ and $L$, the values of $k_{RW}$ and $k_{V}$ are computed via the shooting method. Their dependence on $q/\mu$ is shown in Fig.3 for $D=10,L=2,3,4$. Our results are consistent with those given in [6]. When $q=0$, our results agree with the previously known behavior for Schwarzschild black hole [39]. On the other hand, different from the neutral case, non-zero vector Love numbers are obtained for $D>5$ even when $L=n(D-3)\pm 1$, where $n\in\mathbb{N}^{+}$ [6]. We show a few examples in this case in Fig.4.
    Figure 3: These plots show how vector type Love numbers in $D=10$ behave as we change $q/\mu$ and $L$.
    Figure 4: These plots show that vector Love numbers for RN black hole are generically non-zero in $D>5$ even when $L=n(D-3)\pm 1$, where $n\in\mathbb{N}^{+}$, which is different from Schwarzschild black hole.
@@ 3.4 Effective action of scalar type perturbations
In the $s=0$ sector, the gauge-invariant Love numbers are determined from the $C_{trtr}$ component of Weyl tensor and the temporal component of the Maxwell field $A_{t}$. We find that they are related to $\Psi_{\mathrm{Z}}$ and $\Psi_{S}$ via [39]
@@
We now move on to diagonalizing the scalar perturbations. We make the ansatz
@@ 3.5 Scalar-type Love numbers in diverse dimensions
    In this case, the unambiguous Love number is defined using coefficients $R_{\pm}$ which can be computed analytically by substituting the expansion into the equations. Accordingly, the asymptotic behaviors of $\Psi_{S}$ and $\Psi_{Z}$ near the boundary are given by
@@
    $\Psi_{Z}|_{z\rightarrow 0}$ $=(a_{3}B_{+}+a_{4}B_{-})z^{-\tilde{\ell}}+(a_{3}B_{+}R_{+}+a_{4}B_{-}R_{-})z^{\tilde{\ell}+1}\ln z\ .$ (97)
@@
    $k_{Z}=\frac{a_{3}\lambda_{Z}R_{+}+a_{4}R_{-}}{a_{3}\lambda_{Z}+a_{4}}\ln z=\frac{c_{2}^{2}R_{+}+(\ell-1)(D+\ell-2)R_{-}}{c_{2}^{2}+(\ell-1)(D+\ell-2)}\ln z\ .$ (99)
    Here, $\lambda_{S}=-a_{4}/a_{3}$ and $\lambda_{Z}=-a_{2}/a_{1}$ are the relative amplitudes determined by setting the one of the source term to zero, and $c_{2}$ is defined in (90).
@@
    For $\tilde{\ell}\in\frac{1}{2}+\mathbb{N}$, $R_{\pm}$ takes non-zero real value and the Love numbers exhibit logarithmic running behaviors. We show the results in Fig. 6 for the the case of $D=5$ and $L=3,5,7$.
    Figure 6: Scalar Love numbers for RN black hole in $D=5$ with $L=3,5,7$, where $\ln z$ is omitted.
@@
    $k_{S}=\frac{c_{2}^{2}k_{-}+(\ell-1)(D+\ell-2)k_{+}}{c_{2}^{2}+(\ell-1)(D+\ell-2)},\quad k_{Z}=\frac{c_{2}^{2}k_{+}+(\ell-1)(D+\ell-2)k_{-}}{c_{2}^{2}+(\ell-1)(D+\ell-2)}\ .$ (101)
    The values of $k_{S}$ and $k_{Z}$ are computed via the shooting method for given $D$ and $L$. Examples corresponding to $D=10$ and $L=2,3,4$ are shown in Fig.7. In the neutral limit $q\rightarrow 0$, the numerical results agree with the one in [39].
    Figure 7: Scalar Love numbers for $D=10$ RN black holes with $L=2,3,4$.

### v2

@@ 1 Introduction
Tidal Love numbers describe the deformability of a black hole under the influence of external force. It has been demonstrated that conservative linear tidal Love numbers vanish for several families of asymptotically flat black holes in four dimensions, including Schwarzschild [1], Kerr [2, 3, 4] and Reissner-Nordström (RN) [5, 6, 7] black holes. It was also found that for Kerr black holes, the conservative tidal Love numbers remain vanishing at nonlinear level [8, 9, 10, 11]. These results provide strong support to the no-hair theorem, showing that black holes in four-dimensional general relativity are remarkably rigid objects. On the other hand, Love numbers are generically nonvanishing for charged black holes coupled to charged scalar or fermion [12, 13, 14, 15] and black holes in higher dimensions [16, 17, 18, 19], in modified gravitational theories [20, 21, 22, 23, 24, 25, 26], in the presence of a cosmological constant [27, 28], or in astrophysical environments [29, 30, 31, 32, 33]. Its potential observability by future gravitational wave experiments makes Love numbers useful theoretical diagnostics.
@@
When restricted to $D=4$, our master equations match with results in [43, 44, 45]. In $D>4$, the metric perturbations contain also an intrinsic tensor mode that propagates on its own without mixing with others. This constitutes the simplest perturbation channel. The other components of the metric perturbations inevitably mix with the electromagnetic counterparts. Due to the different methods used in diagonalizing the perturbations, our master equations do not look the same as those in [6]. However, the Love numbers in the tensor and vector sectors do agree with those obtained in [6]. In particular, we confirm that all the Love number vanishes for $4D$ RN black holes [6, 7], a feature similar to the Schwarzschild black hole [1, 2, 40]. In $D>4$, we also recover earlier results for the tensor and vector Love numbers [6]. So far, the scalar type Love numbers of RN black hole have not been studied. In $D$-dimensional spacetimes, the effective multipolar index $\tilde{\ell}$ is related to the original multipolar index $\ell$ by $\tilde{\ell}=\ell/(D-3)$. We recall that in the case of Schwarzschild black hole, when $\tilde{\ell}$ is integer valued, all scalar type Love numbers vanish. Hence, it should be interesting to see if similar features appear also in RN black hole. Our results show that this is indeed the case, which may imply the existence of certain accidental symmetry, as was found for other cases, including $4D$ Schwarzschild black hole [46, 47, 48] and RN black hole [7] and certain higher dimensional black branes [49] or black holes [50].
@@ 3.1 Tensor Love numbers in diverse dimensions
    In this way of parameterization, $\Phi_{\rm resp}$ and $\Phi_{\rm tidal}$ are analytic functions in a neighborhood of $z=0$. The ratio between $A$ and $B$ is fixed, once demanding regularity of the solution at the horizon. Hence the solution takes the form
@@
    The numerical value of $k_{T}$ is determined by matching the solutions near the horizon $(z=1)$ and at infinity $(z=0)$. The regularity at the horizon serves as the key physical boundary condition. This requirement dictates that only the regular solution branch is selected near the horizon. By performing a series expansion in this region, the branch with $\log(1-z)$ is discarded. The regular solution at $z=1$ is solved as a series expansion in powers of (z1)(z-1) up to order $(z-1)^{15}$. We then evaluate the series and its derivative at $z=1-10^{-8}$ which are used as initial values in the numerical integration of the differential equation towards $z=0$. The resulting numerical solution is matched with the asymptotic behavior near $z=0$ (43) to determine the value of $k_{T}$. In Fig.1, we plot both the analytical (solid line) and numerical (points) results for the tensor Love numbers in several cases obeying $2\tilde{\ell}+1\notin\mathbb{N}$. The excellent agreement between the numerical and theoretical values demonstrates the validity of our method.
@@ 3.2 Effective action of vector type perturbations
Unlike the tensor sector (30), the Regge-Wheeler mode in the gravitational perturbation and the magnetic mode in the Maxwell field are both vector-type perturbations, and thus can mix with each other. As a result, the action in the vector sector is more involved than that in the tensor sector (30). In this section, we will show how to diagonalize the vector-type perturbations by introducing an auxiliary field.
Since both the gravitational and Maxwell field perturbations contain vector modes, they are coupled to each other. As a result, the effective Lagrangian in this sector is of the form
@@
The gauge-invariant vector-type Love numbers are encoded in the magnetic components of the Maxwell field strength and linearized Weyl tensor,
@@ 3.3 Vector-type Love numbers in diverse dimensions
In the static case $\omega=0$, solutions to the equations above can be expressed in terms of the Heun function. After imposing regular boundary condition at the outer horizon $x=0$, the solution is
$R_{\mathcal{W}\pm}$ $=$ $\big((1-\sigma)x+1\big)^{-\frac{D-2}{D-3}}\text{HeunG}\Big[\frac{1}{1-\sigma},q_{\pm},\alpha,\beta,1,1,-x\Big]\ ,$
$q_{\pm}$ $=$ $-\frac{\mp c_{\mathcal{W}}+2\big((D-3)^{2}\tilde{\ell}(\tilde{\ell}+1)-D+2\big)+(D-3)(D-1)(1+\sigma)}{2(D-3)^{2}(1-\sigma)}\ ,$
$\alpha$ $=$ $-\frac{D-2}{D-3}-\tilde{\ell},\quad\beta=\tilde{\ell}-\frac{1}{D-3}\ .$ (67)
Had we known the connection for Heun function, the solution above would then be expressed as a linear combination of two other solutions with definite and distinct fall-offs near infinity, from which we read off the Love numbers. However, due to the lacking of the connection formula for Heun function, one still needs to employ numerical method. We thus adopt the same approach as in the previous section which applies to all Fuchsian type equations.
Through a coordinate transformation $x\rightarrow z=\left(\frac{r_{+}}{r}\right)^{D-3}=1/((1-\sigma)x+1)$, we can convert (65) into the standard Fuchsian form. For equations of this type, the indices near the singular point $z=0$ are $\tilde{\ell}$ and $\tilde{\ell}+1$ and the analysis can be divided into the two cases:
@@
    Near spatial infinity ($z\rightarrow 0$) , the asymptotic behaviors of $\Psi_{\mathrm{RW}}$ and $\Psi_{V}$ is then
    $r^{\frac{2-D}{2}}\Psi_{\mathrm{RW}}|_{z\rightarrow 0}$ $=z^{\tilde{\ell}+1}\left(\alpha_{1}B_{+}R_{+}+\alpha_{2}B_{-}R_{-}\right)\ln z+z^{-\tilde{\ell}}(\alpha_{1}B_{+}+\alpha_{2}B_{-})\ ,$ (74)
@@
    $k_{\mathrm{RW}}=\frac{\alpha_{1}B_{+}R_{+}+\alpha_{2}B_{-}R_{-}}{\alpha_{1}B_{+}+\alpha_{2}B_{-}}\ln{z}\ ,\quad k_{V}=\frac{\alpha_{3}B_{+}R_{+}+\alpha_{4}B_{-}R_{-}}{\alpha_{3}B_{+}+\alpha_{4}B_{-}}\ln{z}\ ,$ (76)
    which clearly exhibits gravito-photon mixing. Thus in general, the Love number depends on the ratio of the amplitudes on the horizon. Following [6], when computing the Love number for a specific mode, we set the tidal field of the other mode to zero, thereby fixing the relative amplitude between the two modes. Thus, when calculating $k_{\mathrm{RW}}$ ($k_{V}$), we set the coefficient of the $z^{-\tilde{\ell}}$ term in $\Psi_{V}|_{z\rightarrow 0}$ ($\Psi_{\mathrm{RW}}|_{z\rightarrow 0}$) to zero, fixing the ratio of the amplitudes to be $B_{+}/B_{-}=-\alpha_{4}/\alpha_{3}$ ($B_{+}/B_{-}=-\alpha_{2}/\alpha_{1}$). This allows us to derive $k_{\mathrm{RW}}$ and $k_{V}$ analytically. A special case happens in $D=4$, as one finds $R_{\pm}$ vanishes. We verify that the tidal field $\Phi_{\rm tidal\pm}$ is given by polynomials of degree less than $2\tilde{\ell}+1$ and thus $k_{\mathrm{RW}}=k_{V}=0$. In $D>4$, $R_{\pm}$ is in general nonvanishing. For $D=5$, we list values of the Love numbers (logarithm omitted) for various $\ell$ in Fig.2.
    Figure 2: These plots show how vector type Love numbers in $D=5$ behave as we change $q/\mu$ and $\ell$. $\ln z$ is omitted in the plots.
@@
    $k_{\mathrm{RW}}=\frac{\alpha_{2}\alpha_{3}k_{-}-\alpha_{1}\alpha_{4}k_{+}}{\alpha_{2}\alpha_{3}-\alpha_{1}\alpha_{4}}\ ,\quad k_{V}=\frac{\alpha_{1}\alpha_{4}k_{-}-\alpha_{2}\alpha_{3}k_{+}}{\alpha_{1}\alpha_{4}-\alpha_{2}\alpha_{3}}\ .$ (78)
    For given $D$ and $\ell$, the values of $k_{\mathrm{RW}}$ and $k_{V}$ are computed numerically as before. Their dependence on $q/\mu$ is shown in Fig.3 for $D=10,\ell=2,3,4$. Our results are consistent with those given in [6]. When $q=0$, our results agree with the previously known behavior for Schwarzschild black hole [40]. On the other hand, different from the neutral case, non-zero vector Love numbers are obtained for $D>5$ even when $\ell=n(D-3)\pm 1$, where $n\in\mathbb{N}^{+}$ [6]. We show a few examples in this case in Fig.4.
    Figure 3: These plots show how vector type Love numbers in $D=10$ behave as we change $q/\mu$ and $\ell$.
    Figure 4: These plots show that vector Love numbers for RN black hole are generically non-zero in $D>5$ even when $\ell=n(D-3)\pm 1$, where $n\in\mathbb{N}^{+}$, which is different from Schwarzschild black hole.
@@ 3.4 Effective action of scalar type perturbations
The gauge-invariant scalar type Love numbers are determined from the $C_{trtr}$ component of Weyl tensor and the temporal component of the Maxwell field $A_{t}$. We find that they are related to $\Psi_{\mathrm{Z}}$ and $\Psi_{S}$ via [40]
@@
We now move on to diagonalize the scalar type perturbations. We make the ansatz
@@ 3.5 Scalar-type Love numbers in diverse dimensions
    The unambiguous Love number is usually defined using coefficients $R_{\pm}$ (when it is non-vanishing) which can be computed analytically by substituting the expansion into the equations. Accordingly, the asymptotic behaviors of $\Psi_{S}$ and $\Psi_{\mathrm{Z}}$ near the boundary are given by
@@
    $\Psi_{\mathrm{Z}}|_{z\rightarrow 0}$ $=(a_{3}B_{+}+a_{4}B_{-})z^{-\tilde{\ell}}+(a_{3}B_{+}R_{+}+a_{4}B_{-}R_{-})z^{\tilde{\ell}+1}\ln z\ .$ (98)
@@
    $k_{\mathrm{Z}}=\frac{a_{3}\lambda_{Z}R_{+}+a_{4}R_{-}}{a_{3}\lambda_{Z}+a_{4}}\ln z=\frac{c_{2}^{2}R_{+}+(\ell-1)(D+\ell-2)R_{-}}{c_{2}^{2}+(\ell-1)(D+\ell-2)}\ln z\ .$ (100)
    Here, $\lambda_{S}=-a_{4}/a_{3}$ and $\lambda_{\mathrm{Z}}=-a_{2}/a_{1}$ are the relative amplitudes determined by setting the one of the source term to zero, and $c_{2}$ is defined in (91).
@@
    For $\tilde{\ell}\in\frac{1}{2}+\mathbb{N}$, $R_{\pm}$ takes non-zero real value and the Love numbers exhibit logarithmic running behaviors. We show the results in Fig. 6 for the the case of $D=5$ and $\ell=3,5,7$.
    Figure 6: Scalar Love numbers for RN black hole in $D=5$ with $\ell=3,5,7$, where $\ln z$ is omitted.
@@
    $k_{S}=\frac{c_{2}^{2}k_{-}+(\ell-1)(D+\ell-2)k_{+}}{c_{2}^{2}+(\ell-1)(D+\ell-2)},\quad k_{\mathrm{Z}}=\frac{c_{2}^{2}k_{+}+(\ell-1)(D+\ell-2)k_{-}}{c_{2}^{2}+(\ell-1)(D+\ell-2)}\ .$ (102)
    The values of $k_{S}$ and $k_{\mathrm{Z}}$ are computed via numerical method as before for given $D$ and $\ell$. Examples corresponding to $D=10$ and $\ell=2,3,4$ are shown in Fig.7. In the neutral limit $q\rightarrow 0$, the numerical results agree with the one in [40].
    Figure 7: Scalar Love numbers for $D=10$ RN black holes with $\ell=2,3,4$.